### v2

@@ Abstract
We present a high-dynamic-range vector atomic magnetometer based on applying a fast-rotating magnetic field to a pulsed 87Rb scalar atomic magnetometer. This method enables simultaneous measurements of the total magnetic field and two polar angles relative to the plane of magnetic field rotation. Using two channels in a gradiometer mode, it provides simultaneous measurements of the total field gradient with a sensitivity of 50 $\mathrm{fT/\sqrt{Hz}}$ (1 part per billion), as well as two polar angles with resolutions of 8 $\mathrm{nrad/\sqrt{Hz}}$ at 50 $\mu$T Earth field strength. The noise spectrums of these measurements are flat down to 1 Hz and 0.1 Hz, respectively. Crucially, this approach avoids several metrological difficulties associated with vector magnetometers and gradiometers. We detail the fundamental, systematic, and practical limits of such vector magnetometers. Notably, we provide a comprehensive study of the systematic effects of vector atomic magnetometers. We introduce a new concept of dynamic heading error and investigate several other systematic effects. A unique cosine-altering rotating field modulation is proposed to cancel out these systematics. Additionally, we derive fundamental limits on the sensitivity of such sensors and demonstrate that the vector sensitivity of the sensor can approach its scalar sensitivity while retaining the accuracy and metrological advantages of scalar sensors. This high-dynamic-range vector magnetometer, with ultrahigh resolution and inherent calibration, is suitable for a wide array of applications.
keywords:
@@
1 Introduction
Measurements of magnetic fields have a rich history, starting with the invention of the compass. Among the first and still, most widely used vector sensors are fluxgate magnetometers based on asymmetry in saturation of a soft magnetic material in an applied oscillating magnetic field [1]. Advanced condensed-matter-based approaches involve SQUID magnetometers [2], Hall sensors [3] and magneto-resistive sensors [4]. These sensors are intrinsically vector devices that measure one component of the magnetic field, so three separate sensors are typically required for full vector sensing. For example, the ESA Swarm mission necessitates three-axis fluxgate magnetometers and an Absolute Scalar Magnetometer for calibration [5]. In contrast, atomic magnetometers are inherently scalar devices that gauge the energy gap between spin states in a magnetic field. Additional interactions are required to define the vector axes for a scalar sensor.
Sensitive vector magnetometers operating in Earth’s field have several metrological challenges. For example, a 50 fT magnetic field corresponds to 1 part in $10^{9}$ of Earth’s field. Such relative precision is beyond the capability of most analog voltage measurements. However, frequency measurements can easily achieve much higher relative precision. SQUID magnetometers also can achieve higher fractional resolution using fluxquanta counting techniques [2]. Another challenge is a stable alignment and orthogonality of vector axes directions, which require nano-radian stability. One of the advantages of using external magnetic fields is their relative uniformity that allows the same vector axes orientation for multiple sensors. We use an applied rotating field to create vector sensitivity. For any coil imperfections or non-orthogonality, as long as the coils remain linear, a rotating field always has a unique rotation plane that defines the vector axes in our approach.
@@ 1 Introduction
Several different approaches have created vector sensitivity for atomic sensors. For instance, variometers based on scalar magnetometers [11, 12, 13, 14], magnetic field modulation [15, 16], multi-laser beams to measure multi-spin projections [17, 18, 19, 20], nonlinear magneto-optical rotation (NMOR) magnetometers based on atomic alignment [21, 22], and vector magnetometry based on magneto-optical phenomena [23, 24]. In particular, all-optical vector magnetometers can be achieved by modulation of the light shift, which acts as a pseudo-magnetic field modulation [25]. Lastly, extracting magnetic field vector information by using microwaves as a 3D spatial reference [26]. Some progress in achieving high dynamic range vector magnetometers is listed in Table 1.
Table 1: Vector atomic magnetometers and their performances.
Year Type Tested field ($\mu$T) Total field sensitivity (fT/$\sqrt{\mathrm{Hz}}$) Fractional resolution (1/$\sqrt{\mathrm{Hz}}$) Polar angle sensitivity (nrad/$\sqrt{\mathrm{Hz}}$) Ref
2004 Unshielded SERF magnetometer 50 1,000 (down to 10 Hz) 2×$\times 10^{-8}$ [15]
2006 Variometer based on scalar cesium magnetometer 50 4,000 (estimated) $\times 10^{-8}$ $\times 10^{2}$ (estimated) $\times 10^{2}$ (estimated) [14]
2014 All optical vector NMOR atomic magnetometer 1.5 65 (down to 0.5 Hz) 4×$\times 10^{-8}$ $\times 10^{5}$ only measure one polar angle [25]
2015 Multi-laser beam atomic vector magnetometer based on free spin precession 1 143 (estimated) 1×$\times 10^{-7}$ $4\times 10^{4}$ (estimated) $4\times 10^{4}$ (estimated) [18]
2019 Spin self-sustaining Larmor precession vector magnetometer 0.2 780 (down to 2 Hz) 3.9×$\times 10^{-6}$ 3.5×$\times 10^{3}$ 3.5×$\times 10^{3}$ [27]
2019 Voigt-effect-based vector magnetometer 0.26 400 (down to 2 Hz) 2×$\times 10^{-6}$ 8×$\times 10^{3}$ (down to 10 Hz) 8×$\times 10^{3}$ (down to 10 Hz) [23]
2023 Machine learning assisted NMOR vector magnetometer 0.14 100 (down to 20 Hz) $\times 10^{-7}$ $10^{5}$ (down to 16 Hz) 2×$\times 10^{5}$ (down to 18 Hz) [28]
- FRF vector magnetometer with CARFM modulation 50 50 (down to 1 Hz) 1×$\times 10^{-9}$ 8 (down to 0.1 Hz) 8 (down to 0.1 Hz) This work
@@
We introduce a Fast Rotating Field (FRF) vector magnetometer that can extract three-axis vector information without compromising its scalar resolution. This is achieved by adding a fast rotating magnetic field to compact scalar magnetometers. This approach has enabled the realization of sensitivity as low as several $\mathrm{fT/\sqrt{Hz}}$ in the earth’s ambient environment using multi-pass cells [9, 29].
@@
Our method involves a scalar magnetometer configuration using a pulsed pump laser similar to [9], along with the application of a fast-rotating magnetic field with a rotation rate faster than the spin relaxation rate. This entire sensor can be miniaturized to a very compact size using microelectromechanical systems techniques [30, 31, 32, 33].
@@
We conducted a thorough study on the systematic effects of the magnetometer and developed a unique modulation technique called cosine altering rotating field modulation (CARFM) to cancel out these systematic errors. The magnetometer’s performance is evaluated, and we propose analytical fundamental limits for the sensitivity of the vector free-spin-precession magnetometer based on the Cramer-Rao Lower Bound (CRLB). This is the first comprehensive analysis of the systematic effects of vector atomic magnetometers.
2 Results
2.1 Spin dynamics
@@
The complete description of spin evolution involves density matrix theory [34]. When ground states are distributed according to spin temperature, spin evolution can be effectively described by the Bloch equations [35, 36]. The approximation of rotation matrices is utilized to assess spin evolution in the rotating frame, enabling the derivation of analytical solutions to the time-dependent Bloch equation.
@@
As shown in Fig. 1a, we assume the spins are fully polarized along the negative y-axis. The rotating field initially originates from the x-axis and rotates counterclockwise in the xy-plane (towards the negative z-axis). The spin evolution, without considering spin relaxations, can be expressed as:
$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{\mathbf{e}}_{z}]\cdot\mathcal{R}[\psi,\mathbf{b}]\cdot(-\hat{\mathbf{e}}_{y}),$ (1)
@@
where $\mathcal{R}[\phi,\mathbf{\nu}]$ is a 3D rotation matrix for an anti-clockwise rotation of $\phi$ degrees around the vector $\mathbf{\nu}$. For the applied rotating field in the x-y plane, we have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $\text{Abs}[\phi_{x}-\phi_{y}]=\pi/2$. Therefore, $\mathbf{b}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$. Here, $\gamma$ represents the gyromagnetic ratio, and $\hat{\mathbf{e}}_{x}$, $\hat{\mathbf{e}}_{y}$ and $\hat{\mathbf{e}}_{z}$ are the unit vectors along the x, y and z axes, respectively. We define $\theta=\omega_{m}t$ and $\psi=\gamma t\sqrt{B_{z}-\omega_{m}/\gamma}$. Eventually, the spin projections can be written as:
$P_{x}(t)=$ $\hat{\mathbf{e}}_{x}\cdot\mathbf{P}(t)=\cos\omega_{0}t\sin\omega_{m}t+$ (2)
@@
$P_{z}(t)=$ $\hat{\mathbf{e}}_{z}\cdot\mathbf{P}(t)=-\frac{\gamma B_{m}}{\omega_{0}}\sin\omega_{0}t,$ (3)
where $\omega_{0}=\gamma\sqrt{B_{m}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq.1, we can get
@@
the residual magnetic fields in the x-direction, $b_{x}$, and y-direction, $b_{y}$, are proportional to the amplitudes of the first harmonic, $\sin\omega_{m}t$ and $\cos\omega_{m}t$, respectively, in the phase shift. The conversion factor for converting from the amplitude of the phase shift to the transverse magnetic field is given by ±$\pm\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}/{\gamma B_{m}}$ (Table 2 details the sign dependence). Thus, by analyzing the phase shifts of the Larmor precession signal under the fast rotating field, we can measure both the transverse magnetic fields and the total magnetic field.
b) Four-shot rotating field scheme
@@
Figure 1: (a) Three coils determine the x, y, and z directions. The sensor head sits on a rotating stage with the cell positioned at the stage’s center. The sensor head can freely rotate in the x-z and y-z planes. The angle between the sensor head’s central axis and the z-axis is denoted as $\alpha$. The angle between the sensor head’s axis and the z-axis is labeled as $\beta$. $\theta$ represents the angle between $\mathbf{B_{tot}}$ and $\mathbf{\hat{e}_{z}}$. Notably, WP stands for Wollaston prism, and Q-PD stands for quadrant photodiode. (b) There are four shots with different start phases of $B_{x}$ and $B_{y}$. This four-shot scheme is specifically developed to cancel out the systematic effects, such as Berry’s phase shift, dynamic heading error, probe beam heading error, eddy current, and the systematic caused by the threshold voltage of the MDA. The magnetic field only alters during the peak values (“cosine switch”) to suppress the eddy current caused by the rotating field. The signals during the preparation time are blanked out. (c) Upper: When a magnetic field of -87 nT is applied in the y-direction, the phase shifts of the FID signal acquired in b) are analyzed by the MDA as a function of time. Lower: when the magnetic fields in the x and y directions are roughly compensated. The first harmonic signals are suppressed and the second harmonic signal can be observed.
2.2 Systematics
@@
To achieve high accuracy, systematic errors need to be studied and eliminated. There are several systematics in the FRF vector magnetometer. By generating a power series expansion of Eq. 2 and analyzing the phase shifts, we can identify two phase-shift terms contributed from $B_{m}$. One is Berry’s phase shift, which can be written as
$\tau_{B}(t)=\frac{B_{m}^{2}\omega_{m}t}{2B_{z}^{2}}\approx(1-\cos\theta)\omega_{m}t,$ (5)
@@
where $\theta\approx B_{m}/B_{z}$ is the angle between the total magnetic field and z-axis as shown in Fig. 1a. The Eq. 5 is consistent with Berry’s phase shift described in Ref. [37]. Another phase shift from Eq. 2 is proportional to $\sin(2\omega_{m}t)$, which can be written as
$\tau_{2nd}(t)\approx-\frac{B_{m}^{2}}{4B_{z}^{2}}\sin(2\omega_{m}t).$ (6)
@@
This phenomenon results in a non-zero second harmonic phase shift even in the absence of a residual magnetic field, and when the rotating field is perfectly symmetric in the x-y plane. Asymmetry in the rotating field along the x and y directions also produces a second harmonic signal.
@@
Regarding the probe beam heading error, we define the angle between the probe beam and the x-axis as $\alpha$, as shown in Fig. 1(b). The spin projection along the probe beam can be written as $P_{prob}(t)=P_{x}(t)\cos\alpha+P_{z}(t)\sin\alpha$. By inserting Eqs. 2 and 3, and analyzing the first harmonic phase shift that depends on $\alpha$, we can determine the phase shift caused by the probe beam heading error
$\tau_{prob}(t)\approx-\frac{B_{m}}{B_{z}}\tan\alpha\cdot\sin\omega_{m}t.$ (7)
@@
Converting it into a magnetic field unit, which is equivalent to a fictitious magnetic field $\omega_{m}/\gamma\cdot\tan\alpha$ along the probe beam direction, we named it a probe beam heading error. This error induces a systematic effect where the measured magnetic field depends on the relative angle between the probe beam and $B_{z}$, analogous to the pump beam heading error.
Additionally, there are two extra sources of systematic errors: heading errors and eddy currents. Heading errors include the commonly adopted static heading error, which depends on the sensor’s orientation with respect to the static magnetic field. We introduce the concept of the dynamic heading error, which occurs when the total magnetic field is rotating rather than static. The combined heading error induced by the total magnetic field is the sum of these static and dynamic heading errors. (More details in Sec. 2.3)
@@
Another systematic error is induced by eddy currents. The altering rotating field can generate eddy currents on the electrically conductive aluminum or $\mu$-metal magnetic shield, which in turn generates a magnetic field. The magnitude of the eddy current magnetic field highly depends on how the rotating magnetic field is altered. (More details in Sec. 2.3)
@@
The signs of these systematic errors depend on the rotating field’s phases, while their absolute magnitudes remain the same. The sign of the dynamic heading error ($\tau_{DH}$) depends on the rotating field’s phases, but the sign of the static heading error does not. We introduce a four-shot scheme (shown in Fig. 1b) by switching the phases of the rotating field in the order listed in Table 2. By averaging these four shots, the systematic errors can be canceled out.

### v3

@@ Abstract
We present a vector atomic magnetometer based on applying a fast-rotating magnetic field to a pulsed 87Rb scalar atomic magnetometer. This method enables simultaneous measurements of the total magnetic field and two polar angles relative to the plane of magnetic field rotation. Using two channels in a gradiometer mode, it provides simultaneous measurements of the total field gradient with a sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ (0.7 part per billion), as well as two polar angles with resolutions of 6 $\mathrm{nrad/\sqrt{Hz}}$ at 50 $\mu$T Earth field strength. The noise spectrums of these measurements are flat down to 1 Hz and 0.1 Hz, respectively. Crucially, this approach avoids several metrological difficulties associated with vector magnetometers and gradiometers. We detail the fundamental, systematic, and practical limits of such vector magnetometers. Notably, we provide a comprehensive study of the systematic effects of vector atomic magnetometers. We introduce a new concept of dynamic heading error and investigate several other systematic effects. A unique peak-altering fast rotating field modulation is proposed to cancel out these systematics. Additionally, we derive fundamental limits on the sensitivity of such sensors and demonstrate that the vector sensitivity of the sensor can approach its scalar sensitivity while retaining the accuracy and metrological advantages of scalar sensors. This high-dynamic-range vector magnetometer, with ultrahigh resolution and inherent calibration, is suitable for a wide array of applications.
@@
preprint: APS/123-QEP
I Introduction
The measurement of magnetic fields has a long and rich history, beginning with the invention of the compass. The earliest known compass, the “south-governor” (sinan), dates back over 2000 years. Among the first and still, most widely used vector sensors are fluxgate magnetometers based on asymmetry in saturation of a soft magnetic material in an applied oscillating magnetic field [1]. Advanced condensed-matter-based approaches involve SQUID magnetometers [2], Hall sensors [3] and magneto-resistive sensors [4]. These sensors are intrinsically vector devices that measure one component of the magnetic field, so three separate sensors are typically required for full vector sensing. For example, the ESA Swarm mission necessitates three-axis fluxgate magnetometers and an Absolute Scalar Magnetometer for calibration [5]. In contrast, atomic magnetometers are inherently scalar devices that gauge the energy gap between spin states in a magnetic field, detecting a frequency directly proportional to the field strength. This frequency-based measurement approach allows scalar magnetometers to achieve very high relative precision. Additional interactions are required to define the vector axes for a scalar sensor.
Sensitive vector magnetometers operating in Earth’s field have several metrological challenges. For example, a 50 fT magnetic field corresponds to 1 part in $10^{9}$ of Earth’s field. Such relative precision is beyond the capability of most analog voltage measurements. However, frequency measurements can easily achieve much higher relative precision. SQUID magnetometers also can achieve higher fractional resolution using fluxquanta counting techniques [2]. Another challenge is a stable alignment and orthogonality of vector axes directions, which require nano-radian stability. We use an applied rotating field to create vector sensitivity. For any coil imperfections or non-orthogonality, as long as the coils remain linear, a rotating field always has a unique rotation plane that defines the vector axes in our approach.
@@ I Introduction
Several different approaches have created vector sensitivity for atomic sensors. For instance, variometers based on scalar magnetometers [11, 12, 13, 14], magnetic field modulation [15, 16], multi-laser beams to measure multi-spin projections [17, 18, 19, 20], nonlinear magneto-optical rotation (NMOR) magnetometers based on atomic alignment [21, 22], and vector magnetometry based on magneto-optical phenomena [23, 24]. In particular, all-optical vector magnetometers can be achieved by modulation of the light shift, which acts as a pseudo-magnetic field modulation [25]. Lastly, extracting magnetic field vector information by using microwaves as a 3D spatial reference [26]. While all of these methods measure the amplitudes or frequencies of Larmor precession, frequency measurements can achieve higher fractional precision and are less affected by drifts in cell temperature, pump and probe laser powers, and wavelengths. We focus on measuring the zero-crossings of the Larmor precession signals, which not only provide frequency information but also capture phase shifts and integrate magnetic field modulations, offering more comprehensive insights into the magnetic fields.
@@
We introduce a Fast Rotating Field (FRF) vector magnetometer that can extract three-axis vector information without compromising its scalar resolution. This is achieved by adding a fast rotating magnetic field—at a rotation rate exceeding the spin relaxation rate—to compact scalar magnetometers. Utilizing multipass cells, these scalar magnetometers achieve sensitivities as low as several $\mathrm{fT/\sqrt{Hz}}$ in the earth’s ambient environment [9, 27].
@@
Our method involves a scalar magnetometer configuration using a pulsed pump laser similar to [9], along with the application of a fast-rotating magnetic field. This entire sensor can be miniaturized to a very compact size using microelectromechanical systems techniques [28, 29, 30, 31].
@@
We conducted a thorough study on the systematic effects of the magnetometer and developed a unique modulation technique called peak-altering fast rotating field modulation to cancel out these systematic errors. The magnetometer’s performance is evaluated, and we propose analytical fundamental limits for the sensitivity of the vector free-spin-precession magnetometer based on the Cramer-Rao Lower Bound (CRLB).
II Results
II.1 Spin dynamics
@@
The complete description of spin evolution involves density matrix theory [32]. When ground states are distributed according to spin temperature, spin evolution can be effectively described by the Bloch equations [33, 34].
$\frac{d\mathbf{P}}{dt}=\gamma\mathbf{P}\times\mathbf{B},$ (1)
where $\gamma$ represents the gyromagnetic ratio, $\mathbf{P}$ is spin polarization, and $\mathbf{B}$ is magnetic field. The approximation of rotation matrices is utilized to assess spin evolution in the rotating frame, enabling the derivation of analytical solutions to the time-dependent Bloch equation.
@@
As shown in Fig. 1a, we assume the spins are fully polarized along the negative y-axis. The magnetic field initially points along the z-axis, while a rotating field originates from the x-axis and rotates counterclockwise in the xy-plane. We consider the behavior of polarization in a magnetic field, where it undergoes precession in response to the applied field. This magnetic field is composed of both a static and a rotating component. In the presence of these components, the polarization vector evolves according to specific rotations. The spin evolution, without considering spin relaxations, can be expressed as:
$\mathbf{P}(t)=\mathcal{R}[\theta,\hat{z}]\cdot\mathcal{R}[\psi,\mathbf{B_{tot}}]\cdot(-\hat{y}),$ (2)
@@
where $\mathcal{R}[\phi,\mathbf{v}]$ is a 3D rotation matrix for an anti-clockwise rotation of $\phi$ degrees around the vector $\mathbf{v}$. For the applied rotating field in the xy-plane, we have $B_{x}=B_{m}\sin(\omega_{m}t+\phi_{x})$, $B_{y}=B_{m}\sin(\omega_{m}t+\phi_{y})$, and $|\phi_{x}-\phi_{y}|=\pi/2$. Therefore, $\mathbf{B_{tot}}=(B_{m},0,B_{z}-\omega_{m}/\gamma)$. Here, $\hat{x}$, $\hat{y}$ and $\hat{z}$ are the unit vectors along the x, y and z axes, respectively. We define $\theta=\omega_{m}t$ and $\psi=\gamma t|\mathbf{B_{tot}}|$. Eventually, the spin projections can be written as:
$P_{x}(t)=$ $\hat{x}\cdot\mathbf{P}(t)=\cos\omega_{0}t\sin\omega_{m}t+$ (3)
@@
$P_{z}(t)=$ $\hat{z}\cdot\mathbf{P}(t)=-\frac{\gamma B_{m}}{\omega_{0}}\sin\omega_{0}t,$ (4)
where $\omega_{0}=\gamma|\mathbf{B_{tot}}|$. By inserting residual transverse magnetic fields $b_{x}$ and $b_{y}$ into Eq. 2, we can get
@@
where $|\mathbf{B_{tot}}|=\sqrt{B_{m}^{2}+b_{x}^{2}+b_{y}^{2}+(B_{z}-\omega_{m}/\gamma)^{2}}$. The residual magnetic fields in the x-direction, $b_{x}$, and y-direction, $b_{y}$, are proportional to the amplitudes of the first harmonic, $\sin\omega_{m}t$ and $\cos\omega_{m}t$, respectively, in the phase shift. The conversion factor for converting from the amplitude of the phase shift to the transverse magnetic field is given by ±$\pm\omega_{m}\sqrt{B_{m}^{2}+B_{z}^{2}}/{\gamma B_{m}}$ (Table 1 details the sign dependence). Thus, by analyzing the phase shifts of the Larmor precession signal under the fast rotating field, we can measure both the transverse magnetic fields and the total magnetic field.
The intuitive classical model underlying the operating principle of the FRF vector magnetometer is illustrated in Fig. 1a. The scalar magnitude of the total magnetic field, $|\mathbf{B_{tot}}|$, depends on the angle between the rotating field $\mathbf{B_{m}}$ and the residual magnetic field $\mathbf{B_{0}}$. When a small residual transverse magnetic field $\mathbf{\Delta B}$—comprising $\mathbf{b_{x}}$ and $\mathbf{b_{y}}$—is present, $|\mathbf{B_{tot}}|$ varies with oscillating components at the rotation frequency of $\mathbf{B_{m}}$. The phases and amplitudes of these oscillations are influenced by the angle between $\mathbf{B_{0}}$ and $\mathbf{B_{m}}$. Consequently, the Larmor precession frequency, proportional to $|\mathbf{B_{tot}}|$, also displays oscillating components: a residual field in the x-direction, $b_{x}$, produces an in-phase oscillation, while a residual field in the y-direction induces an out-of-phase component.
This model provides a fundamental description of the magnetometer’s operation. However, for precise measurements, several systematic effects—some even resulting from hyperfine interactions—require in-depth analysis to account for and minimize potential measurement biases.
@@
Figure 1: (a) Three coils determine the x, y, and z directions. The sensor head sits on a rotating stage with the cell positioned at the stage’s center. The sensor head can freely rotate in the xz and yz-planes. The angle between the sensor head’s central axis and the z-axis is denoted as $\alpha$. $\beta$ is the angle by which the pump beam moves away from the position where the pump laser is perpendicular to the magnetic field. $\theta$ represents angle between the axis of rotation and the rotating field vector. Notably, WP stands for Wollaston prism, and Q-PD stands for quadrant photodiode. (b) Trig signal, when high after a rising edge, indicates the preparation phase: the AC heater heats the cell, and the pump laser polarizes the atoms. When low after a falling edge, the pump laser and AC heater are turned off, and the probe beam laser measures the FID signal. There are four shots with different start phases of $B_{x}$ and $B_{y}$.The start phases of the rotating field play a significant role; therefore, we indicate them with vertical dashed lines. This four-shot scheme is specifically developed to cancel out the systematic effects, such as Berry’s phase shift, dynamic heading error, probe beam heading error, eddy current, and the systematic caused by the threshold voltage of the MDA. The magnetic field only alters during the peak values (“Peak-Altering”) to suppress the eddy current caused by the rotating field. The signals during the preparation time are blanked out. (c) Upper: When a magnetic field of -87 nT is applied in the y-direction, the phase shifts of the FID signal acquired in b) are analyzed by the MDA as a function of time. Lower: when the magnetic fields in the x and y directions are roughly compensated. The first harmonic signals are suppressed and the second harmonic signal can be observed.
II.2 Systematics
@@
To achieve high accuracy, systematic errors must be carefully studied and eliminated. In the FRF vector magnetometer, several systematic effects are present. By expanding Eq. 3 into a power series and examining the phase shifts to second order in $B_{m}$, we can identify two phase-shift contributions from $B_{m}$:
[Berry’s phase shift] The Berry phase shift refers to the geometric phase acquired by a spin system during an adiabatic, cyclic evolution [35]. As the magnetic field rotates slowly, the spins adiabatically follow the field. At the end of a complete cycle, when the field returns to its original configuration, the spins acquire a phase shift. If the system undergoes cyclic adiabatic evolution at a constant frequency, the Berry phase accumulates linearly over time. Thus, the Berry phase term as a function of time, from Eq. 3, can be written as
$\phi_{B}(t)=\frac{B_{m}^{2}\omega_{m}t}{2B_{z}^{2}}\approx(1-\cos\theta)\omega_{m}t,$ (6)
@@
where $\theta\approx B_{m}/B_{z}$ is the angle between the total magnetic field and z-axis as shown in Fig. 1a. The Berry’s phase acquired over one complete period, given by Eq. 6 as $2\pi(1-\cos{\theta})$, is consistent with Berry’s phase shift of Nuclear Magnetic Resonance (NMR) described in Ref. [35].
[Second harmonic phase shift] Another phase shift from Eq. 3 is proportional to $\sin(2\omega_{m}t)$, which can be written as
$\phi_{2nd}(t)\approx-\frac{B_{m}^{2}}{4B_{z}^{2}}\sin(2\omega_{m}t).$ (7)
@@
This phenomenon leads to a non-zero second harmonic phase shift, even in the absence of a residual magnetic field and when the rotating field is perfectly symmetric in the xy-plane. Any asymmetry in the rotating field along the x and y-directions also produces a second harmonic signal instead of the first harmonic at $\omega_{m}$, ensuring that it does not affect the measurement of transverse magnetic fields.
[Heading errors] The heading error of atomic magnetometers refers to the dependence of the measured magnetic field values on the orientation of the sensor relative to the magnetic field [36]. The pump beam heading error is well-studied and is primarily caused by nonlinear Zeeman splitting and the difference between Zeeman resonances in the two hyperfine ground states. We set nuclear Landé factor $g_{I}\approx 0$ [37] and the heading error can be well described by $B_{SH}=\mathcal{B_{H}}\sin\beta$, where
$\beta$ is the angle by which the pump beam moves away from the position where the pump laser is perpendicular to the magnetic field. $A_{hf}=h\cdot 3.417$ GHz is the hyperfine structure constant for ground state [38], $h$ is Planck constant. We refer it to as the static heading error (to distinguish it from the dynamic heading error introduced below, we refer to this conventional heading error as static heading error), which depends on the sensor’s orientation with respect to the static magnetic field. We introduce the concept of the dynamic heading error, which occurs when the total magnetic field is rotating rather than static.
We discovered that the dynamic heading error can be explained by the fact that the spin precession plane adiabatically follows the rotation of the total magnetic field. This maintains a constant relative angle between the total magnetic field vector and the spin precession plane, resulting in a stable precession frequency, as the weighted sum of different Zeeman transitions remains consistent. Consequently, the dynamic heading error depends on the initial angle between the spin polarization and the total magnetic field, which remains constant and equal to its initial value. (More details in Sec.  II.3).
To clarify the elimination of the systematic dynamic heading error, we separate the dynamic heading error, $B_{DH}$, into two components: a static part ($B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta$), which is equivalent to the heading error in the absence of the rotating field ($\approx\mathcal{B_{H}}\sin\beta$) when $\theta$ is small, and a rotating-field-phase-dependent part ($B_{PD}$). Notably, the static heading error is independent of the rotating field phases.
@@
There’s an analogous “heading error” effect from the probe beam. This error creates a systematic effect where the measured transverse magnetic field depends on the relative angle between the probe beam and the plane of the rotating field, analogous to the pump beam heading error. Regarding the probe beam heading error, we define the angle that the probe beam rotates away from the x-axis as $\alpha$, with the sensor rotating in the xz-plane., as shown in Fig. 1(b). The spin projection along the probe beam can be written as $P_{prob}(t)=P_{x}(t)\cos\alpha+P_{z}(t)\sin\alpha$. By inserting Eqs. 3 and 4, and analyzing the first harmonic phase shift that depends on $\alpha$, we can determine the phase shift caused by the probe beam heading error
$\phi_{prob}(t)\approx-\frac{B_{m}}{B_{z}}\tan\alpha\cdot\sin\omega_{m}t.$ (9)
@@
Compared to Eq. 5, this is equivalent to transverse magnetic field $b_{x}$. Converting it into a magnetic field unit, it corresponds to a fictitious magnetic field $-\omega_{m}/\gamma\cdot\tan\alpha$ along the x-direction, which we refer to as a probe beam heading error.
@@
[Eddy current] Another systematic error is induced by eddy currents. The altering rotating field can generate eddy currents on the electrically conductive aluminum or $\mu$-metal magnetic shield, which in turn generates a magnetic field. The magnitude of the eddy current magnetic field highly depends on how the rotating magnetic field is altered. (More details in Sec. II.3)
@@
[Sign dependence of systematics] We discovered that the signs of these systematic errors depend on the phases of the rotating field, while their absolute magnitudes remain unchanged. This characteristic offers an opportunity to eliminate these systematics. The sign of the $B_{PD}$ part of dynamic heading error depends on the rotating field’s phases, but the sign of the static heading error does not. We introduce a four-shot scheme (shown in Fig. 1b) by altering the phases of the rotating field in the order listed in Table 1. By averaging these four shots, the systematic errors can be canceled out.